# TurkishMMLU: Measuring Massive Multitask Language Understanding in Turkish

### v1

@@ 3 Dataset
TurkishMMLU is curated using resources from online learning materials for Turkish high school education. In the Turkish educational system, high school education spans four years, and students take the National University Entrance Exams after completing their studies. This exam contains multiple-choice questions covering various subjects from the curricula. To assist students in preparing for these exams, official and commercial exam preparation booklets, video guides, and online practice tests in multiple-choice question-answering format are available. The Turkish Ministry of Education (MEB) has developed an online platform called the Education Information Network (EBA), which aims to provide electronic resources such as lecture notes, videos, tests and solutions, and interactive books to facilitate the learning process for students. This platform111https://ogmmateryal.eba.gov.tr/panel/MSoruDers.aspx contains multiple-choice questions and their solutions that form the basis of our study.
@@ 4 Evaluation Results
Our main evaluation setup is 5-shot in-context learning evaluation, following the prior evaluation setups in recent LLMs (Team et al., 2024b; OpenAI et al., 2024) on English MMLU (Hendrycks et al., 2021). From the development set proposed in §3, we select a fixed set of questions for each subject and include 5 of them in our few-shot prompt, with the question, multiple-choice options, and the answer. We carefully design these prompts to ensure that each question has a different option (in our dataset, the five options are always A, B, C, D, E) as the answer. For evaluation, we report accuracy by using the lm-evaluation-harness framework from EleutherAI (Gao et al., 2023). For open-source models, we perform log-prob based evaluation; for closed-source models we perform greedy decoding and then parse the prediction.
Our second evaluation is a zero-shot evaluation to compare few-shot and zero-shot performance of the models. Additionally, we evaluate LLMs with a 5-shot chain-of-thought (CoT) evaluation. Especially for questions requiring further reasoning and elaboration, such as mathematics, directly giving answers may be a limitation in our main evaluation. Therefore, we evaluate a wide range of models, including closed-source models, with CoT reasoning (Wei et al., 2022). In this setup, we provide CoT solutions for each question in our few-shots for each subject and perform greedy decoding. We put the final answer option at the end of the solution in the prompts, and then parse the predicted answer in the generated solution.
@@ 6 Limitations
While we believe TurkishMMLU will significantly contribute to Turkish NLP and the design of next multilingual LLMs, it does have some limitations. First, TurkishMMLU is focused solely on text-based assessment. Exploring multimodal questions that involve images or audio is left for future work. Second, the dataset covers high school curriculum and university entrance exam questions in a multiple-choice format. However, future efforts should aim to expand Turkish benchmarking datasets to include assessments of generative abilities and more open-ended questions.
Appendix A Leaderboard

### v2

@@ 3 Dataset
TurkishMMLU is curated using resources from online learning materials for Turkish high school education. In the Turkish educational system, high school education spans four years, and students take the National University Entrance Exams after completing their studies. This exam contains multiple-choice questions covering various subjects from the curricula. To assist students in preparing for these exams, official and commercial exam preparation booklets, video guides, and online practice tests in multiple-choice question-answering format are available. The Turkish Ministry of Education (MEB) has developed an online platform called the Education Information Network (EBA), which aims to provide electronic resources such as lecture notes, videos, tests and solutions, and interactive books to facilitate the learning process for students. This platform111https://ogmmateryal.eba.gov.tr/panel/MSoruDers.aspx contains multiple-choice questions and their solutions that form the basis of our study. The questions are driectly sourced from the platform which ensures that each question is crafted and verified by domain experts, guaranteeing correctness and high quality.
We conducted repeated manual reviews, randomly selecting 30 questions from the complete dataset during each review session to assess their formatting and accuracy. Whenever we detected formatting issues—typically due to irregular presentation on the source website—we revised our parsing code to effectively address these edge cases. This process was continued until we ensured consistent formatting accuracy across multiple consecutive rounds of the sampled questions.
@@ 4 Evaluation Results
Our main evaluation setup is 5-shot in-context learning evaluation, following the prior evaluation setups in recent LLMs (Team et al., 2024b; OpenAI et al., 2024) on English MMLU (Hendrycks et al., 2021). From the development set proposed in §3, we select a fixed set of questions for each subject and include 5 of them in our few-shot prompt, with the question, multiple-choice options, and the answer. We carefully design these prompts to ensure that each question has a different option (in our dataset, the five options are always A, B, C, D, E) as the answer. For evaluation, we report accuracy by using the lm-evaluation-harness framework from EleutherAI (Gao et al., 2023). For open-source models, we perform log-prob based evaluation; for closed-source models we perform greedy decoding and then parse the prediction. In order to ensure validity of parsing, we utilize regex patterns based on the few-shot example structure which can be formulated as Question: ...; Choices: A. ... E. ... ; Answer: [A-E] .
Our second evaluation is a zero-shot evaluation to compare few-shot and zero-shot performance of the models. Additionally, we evaluate LLMs with a 5-shot chain-of-thought (CoT) evaluation. Especially for questions requiring further reasoning and elaboration, such as mathematics, directly giving answers may be a limitation in our main evaluation. Therefore, we evaluate a wide range of models, including closed-source models, with CoT reasoning (Wei et al., 2022). In this setup, we provide CoT solutions for each question in our few-shots for each subject and perform greedy decoding. We put the final answer option at the end of the solution in the prompts, and then parse the predicted answer in the generated solution. We formatted the CoT few-shot examples as they would end with the Turkish phrase of The correct choice is [A-E]. This way we can parse the predicted answer easily from the generated solution.
@@ 6 Limitations
While we believe TurkishMMLU will significantly contribute to Turkish NLP and the design of next multilingual LLMs, it does have some limitations. First, TurkishMMLU is focused solely on text-based assessment. Exploring multimodal questions that involve images or audio is left for future work. Second, the dataset covers high school curriculum and university entrance exam questions in a multiple-choice format. However, future efforts should aim to expand Turkish benchmarking datasets to include assessments of generative abilities and more open-ended questions. One other limitation of our study is the potential risk of knowledge leakage, as some large language models (LLMs) may have been pre-trained on datasets that overlap with or are sourced from similar data used in our benchmarks, which could artificially inflate their performance. Although the platform does not display gold answers in its interface, models that encountered these questions during pretraining may still perform better than others. As noted by Pezeshkpour and Hruschka (2024), the order of choices may cause variations in model performance; however, we maintained the original order of the choices in our dataset.
Acknowledgements
This work was funded by Deutsche Forschungsgemeinschaft (project SCHU 2246/14-1).
Appendix A Question Examples
Appendix B Leaderboard
@@
Model | All (macro) | All (micro) | Biology | Physics | Chemistry | Math | Turkish | History | Geography | Philosophy | R&E |
Meta-Llama-3-70B | 66.0 | 63.3 | 65.4 | 59.6 | 53.4 | 30.6 | 58.0 | 77.9 | 78.7 | 89.8 | 81.1 |
Meta-Llama-3-70B-Instruct | 66.0 | 63.7 | 62.9 | 62.8 | 53.5 | 35.1 | 57.9 | 76.6 | 80.7 | 87.2 | 77.4 |
c4ai-command-r-plus | 59.7 | 56.7 | 51.8 | 48.4 | 42.2 | 25.0 | 56.1 | 74.5 | 73.9 | 87.4 | 77.8 |
Mixtral-8x22B-v0.1 | 55.4 | 53.1 | 49.6 | 48.3 | 44.2 | 32.3 | 46.1 | 63.9 | 66.3 | 79.1 | 68.4 |
aya-23-35B | 53.9 | 50.6 | 45.6 | 47.3 | 36.9 | 24.2 | 47.6 | 62.6 | 66.6 | 81.7 | 72.3 |
c4ai-command-r-v01 | 53.0 | 50.2 | 45.8 | 43.5 | 36.1 | 26.0 | 48.1 | 62.4 | 65.7 | 80.4 | 68.8 |
Meta-Llama-3-8B-Instruct | 45.8 | 43.4 | 39.5 | 37.8 | 34.9 | 23.9 | 38.8 | 52.3 | 55.9 | 68.2 | 60.4 |
Meta-Llama-3-8B | 45.4 | 43.4 | 39.1 | 33.9 | 36.1 | 25.1 | 39.3 | 51.3 | 55.7 | 72.0 | 56.5 |
aya-23-8B | 44.3 | 41.8 | 37.7 | 38.6 | 33.4 | 21.2 | 36.8 | 50.8 | 53.4 | 71.6 | 55.2 |
gemma-7b | 44.1 | 41.9 | 36.4 | 37.8 | 34.4 | 25.6 | 36.5 | 47.9 | 54.0 | 69.0 | 55.2 |
Mixtral-8x7B-Instruct-v0.1 | 43.9 | 41.9 | 39.7 | 37.6 | 33.1 | 26.9 | 36.9 | 47.7 | 53.4 | 62.7 | 56.7 |
aya-101 | 40.0 | 37.2 | 30.6 | 30.1 | 29.3 | 19.7 | 37.8 | 44.8 | 45.0 | 68.4 | 54.2 |
Trendyol-LLM-7b-chat-v1.0 | 37.2 | 35.9 | 33.4 | 35.0 | 30.7 | 22.5 | 36.5 | 41.8 | 42.5 | 51.2 | 41.5 |
Trendyol-LLM-7b-chat-dpo-v1.0 | 34.4 | 33.4 | 31.2 | 32.1 | 29.2 | 20.5 | 34.2 | 39.1 | 40.3 | 47.6 | 35.7 |
Mistral-7B-v0.1 | 32.8 | 32.1 | 32.5 | 31.9 | 31.4 | 23.7 | 28.6 | 34.3 | 36.8 | 43.1 | 32.7 |
Mistral-7B-v0.2-hf | 32.5 | 31.9 | 34.3 | 35.7 | 30.7 | 21.2 | 29.1 | 33.6 | 37.6 | 39.5 | 31.0 |
mt0-xxl | 32.4 | 30.2 | 30.0 | 34.5 | 26.8 | 20.2 | 19.1 | 28.6 | 38.9 | 42.8 | 50.7 |
Mistral-7B-Instruct-v0.2 | 32.0 | 31.4 | 31.1 | 32.4 | 30.4 | 23.5 | 32.5 | 29.7 | 36.8 | 43.1 | 28.3 |
Trendyol-LLM-7b-base-v1.0 | 31.1 | 30.0 | 28.7 | 26.2 | 27.2 | 20.2 | 28.3 | 34.7 | 33.9 | 45.9 | 35.3 |
gemma-7b-it | 26.9 | 26.4 | 25.4 | 26.2 | 25.7 | 23.3 | 23.7 | 28.9 | 27.2 | 33.4 | 27.9 |
mt0-xl | 26.4 | 24.7 | 23.8 | 23.1 | 22.7 | 19.2 | 22.0 | 26.8 | 22.7 | 40.7 | 36.5 |
Phi-3-mini-4k-instruct | 25.7 | 25.8 | 26.9 | 28.6 | 27.7 | 24.7 | 23.2 | 25.5 | 25.4 | 30.3 | 18.9 |
Llama-2-13b-chat-hf | 24.6 | 24.0 | 23.4 | 31.3 | 23.2 | 20.4 | 22.5 | 26.2 | 22.8 | 32.6 | 19.5 |
Llama-2-13b-hf | 24.5 | 24.1 | 25.4 | 25.5 | 23.0 | 21.4 | 22.3 | 24.7 | 24.4 | 29.0 | 24.8 |
gemma-2b | 23.5 | 23.2 | 23.4 | 30.6 | 21.9 | 21.4 | 20.8 | 25.1 | 23.1 | 25.3 | 19.7 |
gemma-2b-it | 23.0 | 23.0 | 23.2 | 21.3 | 23.0 | 22.3 | 22.9 | 23.8 | 22.6 | 27.9 | 20.1 |
mt5-base | 22.4 | 22.1 | 24.0 | 30.1 | 21.9 | 18.6 | 19.1 | 23.1 | 23.1 | 21.1 | 20.5 |
Llama-2-7b-chat-hf | 22.1 | 21.9 | 19.4 | 24.4 | 22.9 | 21.4 | 21.2 | 22.5 | 22.3 | 24.3 | 20.3 |
Llama-2-7b-hf | 22.0 | 22.0 | 21.3 | 21.8 | 23.6 | 21.4 | 22.2 | 21.7 | 21.4 | 24.3 | 20.3 |
mt5-xl | 21.2 | 21.1 | 22.0 | 19.0 | 23.1 | 20.1 | 20.7 | 20.7 | 20.2 | 19.9 | 25.1 |
mt0-large | 20.4 | 20.2 | 17.7 | 15.9 | 22.7 | 19.9 | 20.6 | 19.9 | 19.1 | 23.0 | 25.1 |
mt5-xxl | 20.3 | 20.4 | 22.6 | 21.3 | 21.1 | 20.8 | 20.4 | 19.6 | 19.0 | 20.9 | 17.2 |
mt0-base | 20.3 | 20.6 | 18.1 | 17.0 | 23.2 | 21.4 | 20.7 | 20.4 | 21.5 | 21.1 | 18.9 |
kanarya-2b | 19.3 | 19.5 | 20.5 | 18.0 | 17.6 | 20.8 | 19.3 | 19.1 | 20.5 | 20.4 | 17.7 |
mt5-large | 16.9 | 17.0 | 13.9 | 13.6 | 15.3 | 19.1 | 20.0 | 17.4 | 18.2 | 18.0 | 16.8 |